\begin{thebibliography}{0}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi
\expandafter\ifx\csname bibnamefont\endcsname\relax
  \def\bibnamefont#1{#1}\fi
\expandafter\ifx\csname bibfnamefont\endcsname\relax
  \def\bibfnamefont#1{#1}\fi
\expandafter\ifx\csname citenamefont\endcsname\relax
  \def\citenamefont#1{#1}\fi
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\providecommand{\bibinfo}[2]{#2}
\providecommand{\eprint}[2][]{\url{#2}}

\end{thebibliography}


\begin{thebibliography}{10}

\bibitem{AdaTet-98}
{\sc R.~Adami and A.~Teta}, {\em On the {Aharonov-Bohm} effect}, Lett.
  Math.Phys, 43 (1998), pp.~43--53.

\bibitem{AglGriJacPiSem-96}
{\sc U.~Aglietti, L.~Griguolo, R.~Jackiw, S.-Y. Pi, and D.~Seminara}, {\em
  Anyons and chiral solitons on a line}, Phys. Rev. Lett., 77 (1996),
  pp.~4406--4409.

\bibitem{AndBarJon-99}
{\sc I.~Andri\ifmmode~\acute{c}\else \'{c}\fi{}, V.~Bardek, and L.~Jonke}, {\em
  Quantum fluctuations of the chern-simons theory and dynamical dimensional
  reduction}, Phys. Rev. D, 59 (1999), p.~107702.

\bibitem{AroSchWil-84}
{\sc S.~Arovas, J.~Schrieffer, and F.~Wilczek}, {\em Fractional statistics and
  the quantum {H}all effect}, Phys. Rev. Lett., 53 (1984), pp.~722--723.

\bibitem{BarEtalFev-20}
{\sc H.~Bartolomei, M.~Kumar, R.~Bisognin, A.~Marguerite, J.-M. Berroir,
  E.~Bocquillon, B.~Pla{\c c}ais, A.~Cavanna, Q.~Dong, U.~Gennser, Y.~Jin, and
  G.~F{\`e}ve}, {\em Fractional statistics in anyon collisions}, Science, 368
  (2020), pp.~173--177.

\bibitem{BatGuaOel-06}
{\sc M.~T. Batchelor, X.-W. Guan, and N.~Oelkers}, {\em One-dimensional
  interacting anyon gas: Low-energy properties and haldane exclusion
  statistics}, Phys. Rev. Lett., 96 (2006), p.~210402.

\bibitem{Rabello-96}
{\sc S.~J. Benetton~Rabello}, {\em 1d generalized statistics gas: A gauge
  theory approach}, Phys. Rev. Lett., 76 (1996), pp.~4007--4009.

\bibitem{Bon-etal-Pos-21}
{\sc M.~Bonkhoff, K.~J\"agering, S.~Eggert, A.~Pelster, M.~Thorwart, and
  T.~Posske}, {\em Bosonic continuum theory of one-dimensional lattice anyons},
  Phys. Rev. Lett., 126 (2021), p.~163201.

\bibitem{BriHanVas-92}
{\sc L.~Brink, T.~Hansson, and M.~Vasiliev}, {\em Explicit solution to the
  n-body calogero problem}, Physics Letters B, 286 (1992), pp.~109--111.

\bibitem{Calogero-71}
{\sc F.~Calogero}, {\em Solution of the one-dimensional {$N$}-body problems
  with quadratic and/or inversely quadratic pair potentials}, J. Mathematical
  Phys., 12 (1971), pp.~419--436.

\bibitem{CarGreSan-16}
{\sc L.~Cardarelli, S.~Greschner, and L.~Santos}, {\em Engineering interactions
  and anyon statistics by multicolor lattice-depth modulations}, Phys. Rev. A,
  94 (2016), p.~023615.

\bibitem{ChiEtalCel-22}
{\sc C.~S. Chisholm, A.~Fr\"olian, E.~Neri, R.~Ramos, L.~Tarruell, and
  A.~Celi}, {\em Encoding a one-dimensional topological gauge theory in a
  {Raman-coupled Bose-Einstein} condensate}, Phys. Rev. Res., 4 (2022),
  p.~043088.

\bibitem{ClaEtalChi-18}
{\sc L.~W. Clark, B.~M. Anderson, L.~Feng, A.~Gaj, K.~Levin, and C.~Chin}, {\em
  Observation of density-dependent gauge fields in a bose-einstein condensate
  based on micromotion control in a shaken two-dimensional lattice}, Phys. Rev.
  Lett., 121 (2018), p.~030402.

\bibitem{CorDubLunRou-19}
{\sc M.~Correggi, R.~Duboscq, N.~Rougerie, and D.~Lundholm}, {\em Vortex
  patterns in the almost-bosonic anyon gas}, Europhysics Letters, 126 (2019),
  p.~20005.

\bibitem{CorFer-21}
{\sc M.~Correggi and D.~Fermi}, {\em Magnetic perturbations of anyonic and
  {Aharonov-Bohm Schr\"odinger} operators}, J. Math. Phys., 62 (2021),
  p.~032101.

\bibitem{CorOdd-18}
{\sc M.~Correggi and L.~Oddis}, {\em Hamiltonians for two-anyon systems}, Rend.
  Mat. Appl., 39 (2018), pp.~277--292.

\bibitem{VeiOuv-94}
{\sc A.~Dasni\`eres~de Veigy and S.~Ouvry}, {\em Equation of state of an anyon
  gas in a strong magnetic field}, Phys. Rev. Lett., 72 (1994), pp.~600--603.

\bibitem{VeiOuv-95}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em {One-dimensional
  statistical mechanics for identical particles: the {C}alogero and anyon
  cases}}, Modern Physics Letters B, 9 (1995), pp.~271--283.

\bibitem{EdmEtalOhb-13}
{\sc M.~J. Edmonds, M.~Valiente, G.~Juzeli\ifmmode~\bar{u}\else \={u}\fi{}nas,
  L.~Santos, and P.~\"Ohberg}, {\em Simulating an interacting gauge theory with
  ultracold bose gases}, Phys. Rev. Lett., 110 (2013), p.~085301.

\bibitem{Forte-92}
{\sc S.~Forte}, {\em Quantum mechanics and field theory with fractional spin
  and statistics}, Rev. Mod. Phys., 64 (1992), pp.~193--236.

\bibitem{FreMoo-21}
{\sc L.~Fresta and P.~Moosavi}, {\em Approaching off-diagonal long-range order
  for $1+1$-dimensional relativistic anyons}, Phys. Rev. B, 103 (2021),
  p.~085140.

\bibitem{Froehlich-88}
{\sc J.~Fr\"ohlich}, {\em Statistics of fields, the {Yang-Baxter} equation, and
  the theory of knots and links}, in Nonperturbative Quantum Field Theory,
  G.~?t~Hooft~et al., ed., Les Houches - Ecole d'Ete de Physique Theorique,
  1988, pp.~71--100.

\bibitem{Froehlich-90}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Quantum statistics
  and locality}, in Proceedings of the Gibbs Symposium, Les Houches - Ecole
  d'Ete de Physique Theorique, 1990, pp.~89--142.

\bibitem{FroEtalTar-22}
{\sc A.~Fr\"olian, C.~S. Chisholm, E.~Neri, C.~R. Cabrera, R.~Ramos, A.~Celi,
  and L.~Tarruell}, {\em Realizing a 1d topological gauge theory in an
  optically dressed {BEC}}, Nature, 608 (2022), pp.~293--297.

\bibitem{Girardeau-60}
{\sc M.~Girardeau}, {\em Relationship between systems of impenetrable bosons
  and fermions in one dimension}, J. Mathematical Phys., 1 (1960),
  pp.~516--523.

\bibitem{Girardeau-06}
{\sc M.~D. Girardeau}, {\em Anyon-fermion mapping and applications to ultracold
  gases in tight waveguides}, Phys. Rev. Lett., 97 (2006), p.~100402.

\bibitem{Goerbig-09}
{\sc M.~O. Goerbig}, {\em Quantum {H}all effects}.
\newblock arXiv:0909.1998, 2009.

\bibitem{GolMenSha-81}
{\sc G.~Goldin, R.~Menikoff, and D.~Sharp}, {\em {Representations of a local
  current algebra in nonsimply connected space and the Aharonov-Bohm effect}},
  J. Math. Phys., 22 (1981), p.~1664.

\bibitem{GreSan-15}
{\sc S.~Greschner and L.~Santos}, {\em Anyon {H}ubbard model in one-dimensional
  optical lattices}, Phys. Rev. Lett., 115 (2015), p.~053002.

\bibitem{Halperin-84}
{\sc B.~I. Halperin}, {\em Statistics of quasiparticles and the hierarchy of
  fractional quantized {H}all states}, Phys. Rev. Lett., 52 (1984),
  pp.~1583--1586.

\bibitem{HanLeiMyr-92}
{\sc T.~Hansson, J.~Leinaas, and J.~Myrheim}, {\em Dimensional reduction in
  anyon systems}, Nuclear Physics B, 384 (1992), pp.~559--580.

\bibitem{HofLapTid-08}
{\sc M.~Hoffmann-Ostenhof, T.~Hoffmann-Ostenhof, A.~Laptev, and J.~Tidblom},
  {\em {Many-particle Hardy Inequalities}}, J. London Math. Soc., 77 (2008),
  pp.~99--114.

\bibitem{IenLec-92}
{\sc R.~Iengo and K.~Lechner}, {\em Anyon quantum mechanics and
  {C}hern-{S}imons theory}, Phys. Rep., 213 (1992), pp.~179--269.

\bibitem{Jain-07}
{\sc J.~K. Jain}, {\em {Composite fermions}}, Cambridge University Press, 2007.

\bibitem{Khare-05}
{\sc A.~Khare}, {\em {Fractional Statistics and Quantum Theory}}, World
  Scientific, Singapore, 2nd~ed., 2005.

\bibitem{Kundu-99}
{\sc A.~Kundu}, {\em Exact solution of double $\ensuremath{\delta}$ function
  {B}ose gas through an interacting anyon gas}, Phys. Rev. Lett., 83 (1999),
  pp.~1275--1278.

\bibitem{LamLunRou-22}
{\sc G.~Lambert, D.~Lundholm, and N.~Rougerie}, {\em {On quantum statistics
  transmutation via magnetic flux attachment}}.
\newblock arXiv:2201.03518, 2020.

\bibitem{LarLun-16}
{\sc S.~Larson and D.~Lundholm}, {\em {Exclusion bounds for extended anyons}},
  Archive for Rational Mechanics and Analysis, 227 (2018), pp.~309--365.

\bibitem{Laughlin-99}
{\sc R.~B. Laughlin}, {\em Nobel lecture: Fractional quantization}, Rev. Mod.
  Phys., 71 (1999), pp.~863--874.

\bibitem{LeiMyr-77}
{\sc J.~M. {Leinaas} and J.~{Myrheim}}, {\em {On the theory of identical
  particles}}, Nuovo Cimento B Serie, 37 (1977), pp.~1--23.

\bibitem{LeiMyr-88}
{\sc J.~M. Leinaas and J.~Myrheim}, {\em Intermediate statistics for vortices
  in superfluid films}, Phys. Rev. B, 37 (1988), pp.~9286--9291.

\bibitem{Lerda-92}
{\sc A.~Lerda}, {\em {Anyons}}, Springer-Verlag, Berlin--Heidelberg, 1992.

\bibitem{LiBha-94}
{\sc S.~Li and R.~K. Bhaduri}, {\em Mapping of the {S}utherland hamiltonian to
  anyons on a ring}.
\newblock arXiv:cond-mat/9404068, 1994.

\bibitem{Lieb-63}
{\sc E.~H. Lieb}, {\em Exact analysis of an interacting {B}ose gas. {II}. {T}he
  excitation spectrum}, Phys. Rev. (2), 130 (1963), pp.~1616--1624.

\bibitem{LieLin-63}
{\sc E.~H. Lieb and W.~Liniger}, {\em Exact analysis of an interacting {B}ose
  gas. {I}. {T}he general solution and the ground state}, Phys. Rev. (2), 130
  (1963), pp.~1605--1616.

\bibitem{LunQva-20}
{\sc D.~Lundholm and V.~Qvarfordt}, {\em {Exchange and exclusion in the
  non-abelian anyon gas}}.
\newblock arXiv:2009.12709, 2020.

\bibitem{LunRou-16}
{\sc D.~Lundholm and N.~Rougerie}, {\em {Emergence of fractional statistics for
  tracer particles in a Laughlin liquid}}, Phys. Rev. Lett., 116 (2016),
  p.~170401.

\bibitem{LunSol-13a}
{\sc D.~Lundholm and J.~P. Solovej}, {\em {Hardy and Lieb-Thirring inequalities
  for anyons}}, Comm. Math. Phys., 322 (2013), pp.~883--908.

\bibitem{LunSol-13b}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Local exclusion
  principle for identical particles obeying intermediate and fractional
  statistics}, Phys. Rev. A, 88 (2013), p.~062106.

\bibitem{LunSol-14}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em {Local exclusion and
  Lieb-Thirring inequalities for intermediate and fractional statistics}}, Ann.
  Henri Poincar\'e, 15 (2014), pp.~1061--1107.

\bibitem{MinVig-22}
{\sc A.~Minguzzi and P.~Vignolo}, {\em Strongly interacting trapped
  one-dimensional quantum gases: Exact solution}, AVS Quantum Sci., 4 (2022),
  p.~027102.

\bibitem{MisEtalZin-22}
{\sc S.~Mistakidis, A.~Volosniev, R.~Barfknecht, T.~Fogarty, T.~Busch,
  A.~Foerster, P.~Schmelcher, and N.~Zinner}, {\em Cold atoms in low dimensions
  -- a laboratory for quantum dynamics}.
\newblock arXiv:2202.11071, 2022.

\bibitem{Moosavi-22}
{\sc P.~Moosavi}, {\em Exact {Dirac-Bogoliubov-de Gennes} dynamics for
  inhomogeneous quantum liquids}.
\newblock arXiv:2208.14467, 2022.

\bibitem{Myrheim-99}
{\sc J.~Myrheim}, {\em Anyons}, in Topological aspects of low dimensional
  systems, A.~Comtet, T.~Jolic{\oe}ur, S.~Ouvry, and F.~David, eds., vol.~69 of
  Les Houches - Ecole d'Ete de Physique Theorique, 1999, pp.~265--413.

\bibitem{NakEtalMan-20}
{\sc J.~Nakamura, S.~Liang, G.~C. Gardner, and M.~J. Manfra}, {\em Direct
  observation of anyonic braiding statistics}, Nature Physics, 16 (2020),
  pp.~931--936.

\bibitem{Ouvry-01}
{\sc S.~Ouvry}, {\em {On the relation between the anyon and {C}alogero
  models}}, Physics Letters B, 510 (2001), pp.~335--340.

\bibitem{Ouvry-07}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em {Anyons and lowest
  Landau level anyons}}, S\'eminaire Poincar\'e, 11 (2007), pp.~77--107.

\bibitem{OuvPol-09}
{\sc S.~Ouvry and A.~Polychronakos}, {\em {The Lieb--Liniger model in the
  infinite coupling constant limit}}, Journal of Physics A: Mathematical and
  Theoretical, 42 (2009), p.~275302.

\bibitem{OuvPol-18}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em {Mapping the
  {C}alogero model on the anyon model}}, Nuclear Physics B, 936 (2018),
  pp.~189--205.

\bibitem{Polychronakos-89}
{\sc A.~Polychronakos}, {\em Non-relativistic bosonization and fractional
  statistics}, Nuclear Physics B, 324 (1989), pp.~597--622.

\bibitem{PosThoTra-17}
{\sc T.~Posske, B.~Trauzettel, and M.~Thorwart}, {\em Second quantization of
  leinaas-myrheim anyons in one dimension and their relation to the
  lieb-liniger model}, Phys. Rev. B, 96 (2017), p.~195422.

\bibitem{Rougerie-hdr}
{\sc N.~Rougerie}, {\em Some contributions to many-body quantum mathematics}.
\newblock arXiv:1607.03833, 2016.
\newblock habilitation thesis.

\bibitem{RouYan-23b}
{\sc N.~Rougerie and Q.~Yang}, {\em {Dimensional reduction for a system of 2D
  anyons}}.
\newblock arXiv:2305.06670, 2023.

\bibitem{Sen-94}
{\sc D.~Sen}, {\em Dimensional reduction of two-dimensional anyons to a
  one-dimensional interacting {Bose} gas}.
\newblock arXiv:cond-mat/9412013, 1994.

\bibitem{Sutherland-71}
{\sc B.~{Sutherland}}, {\em {Quantum Many-Body Problem in One Dimension: Ground
  State}}, J. Mathematical Phys., 12 (1971), pp.~246--250.

\bibitem{Sutherland-71b}
{\sc B.~Sutherland}, {\em Quantum many-body problem in one dimension:
  Thermodynamics}, J. Math. Phys., 12 (1971), pp.~251--256.

\bibitem{TanEggPel-15}
{\sc G.~Tang, S.~Eggert, and A.~Pelster}, {\em Ground-state properties of
  anyons in a one-dimensional lattice}, New Journal of Physics, 17 (2015),
  p.~123016.

\bibitem{ValWesOhb-20}
{\sc G.~Valent\'{\i}-Rojas, N.~Westerberg, and P.~\"Ohberg}, {\em Synthetic
  flux attachment}, Phys. Rev. Research, 2 (2020), p.~033453.

\bibitem{Vathsan-98}
{\sc R.~Vathsan}, {\em Reduction of anyons to one dimension and
  {Calogero-Sutherland}-type models}, International Journal of Modern Physics
  A, 13 (1998), pp.~4123--4132.

\bibitem{Wilczek-82a}
{\sc F.~Wilczek}, {\em {Magnetic flux, angular momentum, and statistics}},
  Phys. Rev. Lett., 48 (1982), p.~1144.

\bibitem{Wilczek-90}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em {Fractional
  Statistics and Anyon Superconductivity}}, World Scientific, Singapore, 1990.

\bibitem{YakEtal-19}
{\sc E.~Yakaboylu, A.~Ghazaryan, D.~Lundholm, N.~Rougerie, M.~Lemeshko, and
  R.~Seiringer}, {\em Quantum impurity model for anyons}, Phys. Rev. B, 102
  (2020), p.~144109.

\bibitem{YakLem-18}
{\sc E.~Yakaboylu and M.~Lemeshko}, {\em Anyonic statistics of quantum
  impurities in two dimensions}, Phys. Rev. B, 98 (2018), p.~045402.

\bibitem{YaoZhaChi-21}
{\sc K.~X. Yao, Z.~Zhang, and C.~Chin}, {\em Domain-wall dynamics in
  {Bose--Einstein} condensates with synthetic gauge fields}, Nature, 602
  (2021), pp.~68--72.

\end{thebibliography}

\documentclass[aps,pra,twocolumn,showpacs,amsmath,amssymb,superscriptaddress]{revtex4}
\usepackage{amsthm,amsmath,amsfonts,amssymb,amsxtra,appendix,bookmark,dsfont,latexsym,stix}

\makeatletter
\usepackage{hyperref}
\usepackage{delarray,color}
\usepackage{euscript}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{remark}

\def\geqslant{\ge}
\def\leqslant{\le}
\def\bq{\begin{eqnarray}}
	\def\eq{\end{eqnarray}}
\def\bqq{\begin{eqnarray*}}
	\def\eqq{\end{eqnarray*}}

\def\eps{\varepsilon}

\newcommand{\norm}[1]{\left\lVert #1 \right\rVert}

\newcommand\1{{\ensuremath {\mathds 1} }}

\newcommand{\im}{\mathrm{i}}

\renewcommand{\epsilon}{\varepsilon}
\newcommand{\bx}{\mathbf{x}}

\def\R {\mathbb{R}}

\def\N {\mathcal{N}}

\def\R {\mathbb{R}}

\def\N {\mathbb{N}}

\def\d{{\rm d}}

\def\b{|}

\renewcommand{\leq}{\leqslant}
\renewcommand{\geq}{\geqslant}

\newcommand{\bA}{\mathbf{A}}

\newcommand{\bp}{\mathbf{p}}

\usepackage{color}

\newcommand{\Htwo}{H^{\rm 2D}_{\varepsilon}}
\newcommand{\Hone}{H^{\rm 1D}}
\newcommand{\HHO}{H^{\rm HO}_\eps}
\newcommand{\Done}{\mathcal{D}^{\rm 1D}}
\newcommand{\ltwo}{\lambda^{\rm 2D}}
\newcommand{\lone}{\lambda^{\rm 1D}}

\newcommand{\St}{\widetilde{S}}

\date{September, 2023}

\begin{document} 

\title{Anyons in a tight wave-guide and the Tonks-Girardeau gas}

\author{Nicolas Rougerie}
\affiliation{Ecole Normale Sup\'erieure de Lyon \& CNRS, UMPA (UMR 5669)}
\email{nicolas.rougerie@ens-lyon.fr}

\author{Qiyun Yang}
\affiliation{Ecole Normale Sup\'erieure de Lyon, UMPA (UMR 5669)}
\email{qiyun.yang@ens-lyon.fr}

\begin{abstract}
We consider a many-body system of 2D anyons, free quantum particles with general statistics parameter $\alpha \in ]0,2[$. In the magnetic gauge picture they are described as bosons attached to Aharonov-Bohm fluxes of intensity $2\pi \alpha$, generating long-range magnetic forces. A dimensional reduction to 1D is obtained by imposing a strongly anisotropic trapping potential. This freezes the motion in the direction of strong trapping, leading to 1D physics along the weak direction. The latter is governed to leading order by the Tonks-Girardeau model of impenetrable bosons, independently of $\alpha$.  
\end{abstract}

\pacs{05.30.Pr, 03.75.Hh,73.43.-f,03.65.-w}

\maketitle

Quasi-particle excitations of many-body systems confined to reduced dimensionalities are not in principle constrained by the symmetry dichotomy which sorts all fundamental particles into bosons and fermions~\cite{LeiMyr-77,GolMenSha-81,Wilczek-82a,Myrheim-99}. In 2D, many-body quantum wave-functions may be classified by the phase picked upon exchanging/braiding two particles. It is of the form $e^{\im \pi \alpha}$ for $\alpha\in [0,2[$, the standard cases of bosons and fermions being recovered for $\alpha = 0,1$ respectively. Equivalently one may think of these so-called anyons in terms of standard bosons (or fermions), coupled to infinitely thin magnetic flux tubes of strength $2\pi \alpha$ (or $2\pi (\alpha-1)$). This point of view is referred to as the magnetic gauge picture~\cite{Froehlich-88,Froehlich-90,Lerda-92,Wilczek-90,IenLec-92,Forte-92,Khare-05,Ouvry-07}. In 1D there does not seem to be a unique agreed-upon model for anyonic exchange statistics. Depending on how one proceeds to quantization, they have historically been described~\cite{LeiMyr-77,LeiMyr-88,Polychronakos-89,PosThoTra-17,Ouvry-07} as ordinary particles with either contact interactions (Lieb-Liniger model) or inverse-square interactions (Calogero-Sutherland model) but other formalisms exist~\cite{Girardeau-06,AndBarJon-99,FreMoo-21,Moosavi-22}. In particular, the chiral BF/Kundu model~\cite{Rabello-96,AglGriJacPiSem-96,Kundu-99,BatGuaOel-06} and the anyon-Hubbard model~\cite{GreSan-15,TanEggPel-15,Bon-etal-Pos-21} have attracted attention recently. Our main purpose here is to ask which one, if any, of the different theoretically possible descriptions of 1D anyons, is singled out as the dimensional reduction of the 2D theory.   

The main candidates for real-world implementation of 2D anyon statistics remain the charge carriers of fractional quantum Hall systems~\cite{Halperin-84,AroSchWil-84,LunRou-16,LamLunRou-22} or their counterparts e.g. in cold atom emulations~\cite{EdmEtalOhb-13,ClaEtalChi-18,ValWesOhb-20,YakEtal-19,YakLem-18,CorDubLunRou-19}. See~\cite{Goerbig-09,Laughlin-99,Jain-07,Rougerie-hdr} for reviews and~\cite{BarEtalFev-20,NakEtalMan-20} for experimental evidence. Said charge carriers are 2D objects described in the bulk via the usual, aforementioned, anyon model~\cite{LunRou-16,LamLunRou-22}. Much of fractional quantum Hall physics is however probed via the transport of charge carriers along 1D edge channels, which connects to our main question.

In another direction, the coupling of cold atoms to optical fields can lead, in the adiabatic limit, to the effective implementation of density-dependent gauge fields~\cite{EdmEtalOhb-13,ValWesOhb-20,CarGreSan-16}. Key proposals in this direction have recently been experimentally realized~\cite{ClaEtalChi-18,FroEtalTar-22,YaoZhaChi-21}. In particular, signatures of the chiral BF model, connected to 1D Kundu anyons, have been observed~\cite{ChiEtalCel-22,FroEtalTar-22} by generating a magnetic-like \emph{vector potential} proportional to matter density. On the other hand, the magnetic gauge picture of 2D anyons corresponds to to a magnetic-like \emph{field} proportional to matter density.

\bigskip 

Indeed, in this note we explain (full mathematical details will be provided elsewhere) that the magnetic-gauge picture Hamitonian for 2D anyons of statistical parameter $\alpha \in ]0,2[$ converges, in the limit of a tight confinement along one spatial dimension, to the impenetrable boson model of the Tonks-Girardeau gas, soluble by Bose-Fermi mapping~\cite{MinVig-22,Girardeau-60,MisEtalZin-22}. Thus, at leading order, the physics does not depend on $\alpha$ and is given by an extreme case of the Lieb-Liniger~\cite{LieLin-63,Lieb-63} model. The behavior is always essentially fermionic~\cite{OuvPol-09}. 

These results might be interpreted in light of the enhanced effect of interactions in reduced dimensionalities. However, it is remarkable that the long-range magnetic interactions of the original model result in a limiting purely local theory. This finding is consistent with~\cite{HanLeiMyr-92,Sen-94} although our approach differs and seems more systematic. In particular it clarifies the vanishing of the long-range magnetic interaction. A particular, discontinuous, phase is acquired by the 2D wave-functions, gauging the interaction away when particles are aligned.

\section{Models and main result}

Consider a (multi-valued) wave function $\Psi :\left (\R^{2}\right )^{N}\to \mathbb{C}$ with anyonic exchange  behavior, i.e.
\begin{equation}\label{eq:stat}
	\Psi (\bx_{1},...,\bx_{j},...,\bx_{k},...,\bx_{N})=e^{\im\alpha \pi}\Psi(\bx_{1},...,\bx_{k},...,\bx_{j},...,\bx_{N})
\end{equation}
with $\alpha \in \left ]0,1\right]$ (by periodicity and complex conjugation, considering this range is sufficient). 
It is convenient to perform a singular gauge transformation
$$\Psi(\bx_{1},...,\bx_{N})=\prod_{j<k}e^{\im\alpha \phi_{jk}}\Phi(\bx_{1},...,\bx_{N}) \;\:\text{with}\;\:
\phi_{jk}=\mathrm{arg}\frac{\bx_{j}-\bx_{k}}{|\bx_{j}-\bx_{k}|}$$
with $\Phi$ a bosonic wave function, symmetric under particle exchange. We have denoted $\mathrm{arg}(\,.\,)$ the angle of a planar vector with the horizontal axis. Applying this transformation, one finds
$$ \left\langle \Psi | \left(-\im\nabla_{\bx_{j}} \right) ^2 | \Psi \right\rangle = \left\langle \Phi | D_{\bx_j} ^2 | \Phi \right\rangle$$
where the momentum operator for particle $j$ has changed as
\begin{equation}\label{eq:kin op}
-\im\nabla_{\bx_{j}} \rightsquigarrow D_{\bx_j}:=-\im\nabla_{\bx_{j}} +\alpha \bA\left (\bx_{j}\right ) 
\end{equation}
with, denoting $\left (x,y\right )^{\perp}=\left (-y,x\right )\in \R^2$,
\begin{equation}
	\bA\left (\bx_{j}\right ):=\sum_{k\neq j}\frac{\left (\bx_{j}-\bx_{k}\right )^{\perp}}{ \b \bx_{j}-\bx_{k} \b^{2}}.
	\label{AJP}
\end{equation}
In this picture we have traded the non-trivial exchange symmetry of wave-functions for a density-dependent magnetic field. Particle $j$ sees all the others as carrying an Aharonov-Bohm flux, leading to the magnetic field 
$$
B(\bx_j) = \mathrm{curl}_{\bx_{j}} \bA \left(\bx_{j}\right ) = 2\pi \alpha \sum_{k\neq j} \delta_{\bx_j=\bx_k}
$$
We adopt this point of view throughout the note, using (the Friedrichs extension~\cite{AdaTet-98,LunSol-13a,LunSol-13b,LunSol-14,CorOdd-18,CorFer-21} of)
\begin{equation}\label{eq:H2D}
\Htwo := \sum_{j=1} ^N \left(D_{\bx_j} ^2 + V_{\eps} (\bx_j) \right)
\end{equation}
acting on bosonic wave-functions as our starting point, where 
\begin{equation}\label{eq:pot}
V_{\eps} (\bx) = V_{\eps} (x,y) = x^2 + \frac{y^2}{\eps ^2}
\end{equation}
is a convenient way of enforcing 1D behavior along the horizontal axis in the limit $\eps \to 0$. This is arguably a crude description (but, perhaps, also an instructive toy model) if one has a fractional quantum Hall edge in mind. There is however no difficulty in imposing such a potential on emerging anyons in cold atoms systems. The choice of a harmonic trapping is only out of convenience. Our results remain true with different choices, but the harmonic trapping has the virtue of leading to exactly soluble limit models.

We denote 
\begin{equation}\label{eq:ener 2D}
\ltwo_k = \min_{\substack{V_k \subset L^2 \left(\R^{2N}\right) \\ \mathrm{dim} V_k = k}} \, \max_{\Psi \in V_k,\norm{\Psi}_{L^2} = 1} \left\langle\Psi| \Htwo| \Psi  \right\rangle 
\end{equation}
the eigenvalues of~\eqref{eq:H2D}, defined by standard Courrant-Fisher min-max formulae. Let $\Psi_k$ be associated eigenfunctions, i.e. 
$$ \Htwo \, \Psi_k = \ltwo_k \Psi_k.$$
There can be degeneracies, in which case we count eigenvalues with their multiplicities. 

\bigskip

For small $\eps >0$ one expects the motion in the two spatial directions to decouple (which is true only to some extent in this particular case, see below). The motion in the $y$ direction will be frozen in the ground state of the harmonic oscillator 
\begin{equation}\label{eq:HO}
\HHO := -\partial_y ^2 + \frac{y^2}{\eps ^2}. 
\end{equation}
It turns out that the motion in the $x$ direction reduces to the free Hamiltonian 
\begin{equation}\label{eq:H1D}
\Hone := \sum_{j=1} ^N -\partial_{x_j} ^2 + {x_j}^2 
\end{equation}
but acting on the domain (of the Friedrichs extension)
\begin{equation}\label{eq:dom1D}
\Done := \left\{ \psi \in H^2 (\R^N), \psi (x_j = x_k) \equiv 0 \mbox{ for all } j\neq k \right\}.
\end{equation}
This restriction is equivalent to the addition of a delta pair-potential of infinite strength to~\eqref{eq:H1D}. It is well-known that this impenetrable boson model can be mapped to a free fermionic one~\cite{MinVig-22,Girardeau-60,MisEtalZin-22}. In turn, this leads to an exact solution in the particular case above. However, our approach does not rely on this exact solution, and we could in fact have added extra interactions to our model. For simplicity, we do not consider this explicitly.

Let $e_\eps$ and $u_\eps$ be respectively the lowest eigenvalue and eigenfunction of~\eqref{eq:HO}. Let $(\lone_k)_{k \in \N}$ be the eigenvalues of~\eqref{eq:H1D}, with associated eigenfunctions $\psi_k, k = 1,2\ldots$. In the model above, $\lone_k$ is a sum of eigenvalues of the harmonic oscillator $-\partial_x + x^2$ and 
$$ \psi_k = c_k \prod_{i<j} \mathrm{sgn} (x_i-x_j) \underset{i,j}{\det} (v_i (x_j))$$
where $v_i$ are the associated one-particle eigenfunctions and $c_k$ is a normalization constant.

\bigskip

We state our main finding as a theorem. 

\begin{theorem}[\textbf{Dimensional reduction for anyons}]\label{thm:main}\mbox{}\\
For all $k\in \N$, in the limit $\eps \to 0$,  
\begin{equation}\label{eq:ener}
\ltwo_k = N e_\eps + \lone_k + o (1). 
\end{equation}
Moreover, one can choose the 2D and 1D eigenbases $(\Psi_k)_k$ and $(\psi_k)_k$ in such a way that
\begin{equation}\label{eq:func}
\int_{\R^{2N}} \left| \Psi_k - \psi_k (x_1,\ldots,x_N)\prod_{j=1} ^N u_{\eps} (y_j)  \right|^2 \underset{\eps \to 0}{\to} 0  
\end{equation}
\end{theorem}

Although it seems from~\eqref{eq:func} that a standard decoupling between the two space directions takes place, the actual ans\"atze for the eigenfunctions $\Psi_k$ leading to the correct energy are more subtle. Essentially they are of the form
\begin{equation}\label{eq:ansatz}
\psi_k (x_1,\ldots,x_N) \prod_{j=1} ^N u_{\eps} (y_j) \prod_{j<k} e^{- \im \alpha S (\bx_j - \bx_k)} 
\end{equation}
where we denote 
\begin{equation}\label{eq:phase}
S (\bx) = S (x,y) = \arctan\left(\frac{y}{x}\right). 
\end{equation}
The above trial states have the correct bosonic symmetry because $S (\bx) = S (-\bx)$, but they are not of the form ``function of $x$ times function of $y$'' that is more common in dimensional reductions. Note that $S(\bx)$ has a discontinuity along the line $x= 0$, so that it is crucial for~\eqref{eq:ansatz} to be well-defined that $\psi_k (x_1,\ldots,x_N)$ vanishes whenever $x_j=x_k,\, j\neq k$.  

The phase factors $e^{-\im\alpha S (\bx_j - \bx_k)}$ modify the energy dramatically, gauging away the original magnetic interaction (see below). For this effect it is crucial to take advantage of the finite, albeit small, extension of our wave-guide, as shown by the discussion below. 

\section{A case for the Calogero-Sutherland model}

Before we sketch the proof of the above, it is instructive to examine an argument that would rather point in the direction of the Calogero model with inverse square interactions (which is also a proposed model for 1D anyons) as effective description. This will emphasize two things: 

\medskip 

\noindent$\bullet$ That quantization and dimensional reduction do not commute in this particular case. Classical particles with the above magnetic interactions would experience Calogero-like interactions if confined on a line.

\medskip

\noindent$\bullet$ The role of the phase factors $e^{-\im\alpha S (\bx_j - \bx_k)}$ in the main result. Indeed, if one chooses a simpler ansatz of the form 
 $$ \psi_ k =  \phi(x_1,\ldots,x_N) \prod_{j=1} ^N u_{\eps} (y_j),$$
the 1D function $\phi$ indeed experiences a Calogero-type Hamiltonian. 

\medskip

The possible connection between 2D anyons and Calogero-like models have already been pointed out in a similar context~\cite{LiBha-94,Vathsan-98}. It also arises for lowest Landau level anyons~\cite{BriHanVas-92,HanLeiMyr-92,Ouvry-01,Ouvry-07,OuvPol-18,VeiOuv-94,VeiOuv-95} via very different mechanisms. 

Consider $N$ classical particles with magnetic interactions akin to those of~\eqref{eq:H2D}. We constrain them to move on the line $y=0$, like perls on a necklace. Expanding the square in~\eqref{eq:H2D}, the Hamilton function for this system is 
\begin{align*}
H (\bx_1,\ldots,\bx_N&;\bp_1,\ldots,\bp_N) = \sum_{j=1} ^N \left(|\bp_j| ^2 + x_j^2 \right)\\
&+ 2 \alpha \sum_{j\neq k} \bp_j \cdot \frac{(\bx_j - \bx_k)^{\perp}}{|\bx_j-\bx_k|^2}\\
&+\alpha^2 \sum_{j \neq k \neq \ell} \frac{(\bx_j - \bx_k)^{\perp}}{|\bx_j-\bx_k|^2} \cdot \frac{(\bx_j - \bx_\ell)^{\perp}}{|\bx_j-\bx_\ell|^2} \\
&+ \alpha^2 \sum_{j\neq k} \frac{1}{|\bx_j-\bx_k|^2} 
\end{align*}
where $\bp_j = (p_j,0)$ and $\bx_j = (x_j,0)$ are momenta and positions, respectively. The cross-term on the second line is clearly null. The term on the third line is null as well, as follows from grouping terms as in~\cite[Lemma~3.2]{HofLapTid-08} 
$$
\sum_{\mbox{cyclic in } 1,2,3} \frac{(\bx_1 - \bx_2)^{\perp}}{|\bx_1-\bx_2|^2} \cdot \frac{(\bx_1 - \bx_3)^{\perp}}{|\bx_1-\bx_3|^2} = \frac{1}{2\mathcal{R}(\bx_1,\bx_2,\bx_3) ^2}
$$
with $\mathcal{R}(\bx_1,\bx_2,\bx_3)$ the circumradius of the triangle with summits $\bx_1,\bx_2,\bx_3$. This is the radius of the circle on which the three points lie, which is infinite for aligned points. Hence the Hamilton function boils down to 
$$
\sum_{j=1} ^N \left(p_j ^2 + x_j^2 \right) + \alpha^2 \sum_{j\neq k} \frac{1}{(x_j-x_k)^2} 
$$
which, once quantized, gives a Calogero Hamiltonian, albeit not with the expected $\alpha (\alpha - 1)$ coefficient~\cite{BriHanVas-92,LeiMyr-88,Polychronakos-89} in front of the two-body term for particles of statistics parameter $\alpha$. Note that this reduction could in any case not be correct for all $\alpha$ because the 2D anyon energy is periodic in $\alpha$, but the Calogero energy is not~\cite{BriHanVas-92,Calogero-71,Sutherland-71,Sutherland-71b}.

\section{Argument for the main result}

We turn to sketching the main insights of the proof of Theorem~\ref{thm:main}. Turning them into a rigorous mathematical proof is somewhat lengthy, and will be done elsewhere~\cite{RouYan-23b}.

The crucial observation is that for particles close to the line $y=0$, the vector potential of the Aharonov-Bohm fluxes in~\eqref{eq:H2D} can be gauged away. The vector potential 
$$ \bA_0 (\bx) = \frac{\bx^\perp}{|\bx|^2} = \begin{pmatrix} -y \\ x \end{pmatrix} \frac{1}{x^2 + y^2}$$
for a unit Aharonov-Bohm flux at the origin has a non-zero curl and thus cannot be written as the gradient of a regular function globally. But, with $S$ defined as in~\eqref{eq:phase}
$$ 
\nabla S (\bx)= \bA_0 (\bx)  -  \begin{pmatrix} \pi \delta_{x=0}\mathrm{sgn} (y) \\ 0 \end{pmatrix}.
$$
Hence, for any continuous function $\Psi (\bx)$ of finite kinetic energy vanishing on the line $x=0$ 
$$ \int_{\R^2} \left| \left(-\im \nabla_{\bx} + \alpha\bA_0 (\bx)\right) \left(\Psi (\bx) e^{-\im \alpha S (\bx)} \right) \right|^2 d\bx = \int_{\R^2} |\nabla \Psi| ^2 $$
and this will be our model calculation (here performed in the relative coordinate of a particle pair). 

The main point of our argument is the behavior 
$$ 
S (\bx) \simeq_{|y| \ll |x|} \frac{y}{x}.
$$
Indeed if one sets instead
$$\St (\bx):=  \frac{y}{x}$$
one finds
$$ 
\nabla \St (\bx)= \begin{pmatrix} - y/x^2 \\ 1/x \end{pmatrix} \underset{|y|\ll |x|}{\simeq} \bA_0 (\bx)
$$
and more precisely 
\begin{equation}\label{eq:diff phase}
\left| \nabla \St - \bA_0 \right| \leq C \frac{|y|}{x^2}. 
\end{equation}
The singularity around $x=0$ of the right-hand side would have to be tamed if we used $\St$ instead of $S$ in our trial state. Hence the latter choice is actually simpler, and we stick to it in the sequel.

%
%
%
%
Consider now a trial state $\Psi$ for~\eqref{eq:H2D} and write it as 
\begin{align}\label{eq:decouple}
 \Psi (\bx_1,\ldots,\bx_N) &= U_\eps \prod_{j<k} e^{-\alpha \im S(\bx_j - \bx_k)} \Phi\\
 U_\eps (\bx_1,\ldots,\bx_N) &= \prod_{j=1} ^N u_\eps (y_j) \nonumber
\end{align}
with a new, continuous, unknown function $\Phi$ vanishing whenever $x_j = x_k, j\neq k$. A direct calculation yields 
\begin{multline}\label{eq:splitting}
\left\langle \Psi |\Htwo| \Psi \right\rangle = Ne_\eps + \sum_j ^N\int_{\R^{2N}} U_\eps ^2 x_j^2 |\Phi|^2  \\
+ \sum_j ^N\int_{\R^{2N}} U_\eps ^2 \left|\nabla_{\bx_j} \Phi \right|^2  
\end{multline}
and we now seek critical points of this functional of $\Phi$. 
For energy upper bounds it will clearly be favorable for $\Phi$ not to depend on $y_1,\ldots,y_N$ and we then recognize the energy functional corresponding to~\eqref{eq:H1D}.
%

We then need to prove a lower bound of the correct form for the energy of a true eigenstate $\Psi$ of the 2D model. We do not know a priori that the corresponding $\Phi$ vanishes when $x_j=x_k,j\neq k$, but~\eqref{eq:splitting} can in this case be replaced by  
\begin{multline}\label{eq:splitting 2}
\left\langle \Psi |\Htwo| \Psi \right\rangle \geq Ne_\eps + \sum_j ^N\int_{\Lambda_\eta} U_\eps ^2 x_j^2 |\Phi|^2  \\
+ \sum_j ^N\int_{\Lambda_\eta} U_\eps ^2 \left|\nabla_{\bx_j} \Phi \right|^2  
\end{multline}
where 
$$ \Lambda_\eta = \left\{ (\bx_1,\ldots,\bx_N) \in \R^{2N}, |x_j- x_k| > \eta, \forall j\neq k \right\}.$$
Extracting the singular phase is unproblematic on the latter set, and we can then pass to the limit in the above, $\eps \to 0$ and $\eta \to 0,$ obtaining an energy lower bound essentially of the desired form. More precisely, in~\cite[Proposition~4.3]{RouYan-23b} we prove the following:

\begin{lemma}
Let 
\begin{equation}\label{eq:rescale func}
\phi(x_1,y_1,\dots,x_N,y_N) := \Phi(x_1,\sqrt{\varepsilon} y_1,\dots,x_N,\sqrt{\varepsilon}y_N). 
\end{equation}
After possibly extracting a subsequence, 
$$ \phi \underset{\varepsilon \to 0}{\to} \psi_{0}$$
where the limit $\psi_0$ has no dependence on the $y$-coordinates, and satisfies
\begin{multline}\label{proofLower7}
    \liminf_{\varepsilon \to 0} \left(\left\langle \Psi |\Htwo| \Psi \right\rangle - Ne_{\varepsilon} \right)
    \\ \ge \sum\limits_{j=1}^N \left( \int_{\mathbb{R}^N} \left|\partial_{{x}_j}{\psi_{0}}\right|^2 + \int_{\mathbb{R}^N} |x_j| ^2 \left|\psi_{0}\right|^2 \right) .
\end{multline}
\end{lemma}

The main difficulty is now to ensure some form of vanishing around particle encounters for our limit model to indeed be set on~\eqref{eq:dom1D}, i.e. that the 1D function obtained by passing to the limit  has finite kinetic energy over the whole space (including across diagonals). To this end we use the following Hardy-like inequality: For any $\Psi$, with the modified momentum as in~\eqref{eq:kin op}   
\begin{multline}\label{eq:Hardy}
 \sum_{j=1} ^N \left\langle \Psi | \left(D_{\bx_{j}} \right) ^2 | \Psi \right\rangle \geq \\ 
 c_{\alpha,N} \int_{\R^{2N}} \sum_{j<k} \frac{1}{|\bx_j - \bx_k| ^2} |\Psi (\bx_1,\ldots,\bx_N)| ^2
\end{multline}
where the best possible constant $c_{\alpha,N}$ depends only on $\alpha$ and $N$. Such an inequality originates from~\cite{HofLapTid-08,LunSol-13a,LunSol-13b,LarLun-16} (see also~\cite{LunQva-20} for review and generalizations) where it is proved in particular that 

\begin{itemize}
 \item $c_{\alpha,N} \geq c N^{-1}$ with a universal $c>0$ if $\alpha$ is an odd-numerator fraction.
 \item $c_{\alpha,2} > 0$ for any $\alpha \neq 0$.
\end{itemize}

\medskip

We improve these bounds by proving that there exists a universal constant $c'>0$ such that
$$c_{\alpha,N} \geq c' N^{-2} \mbox{ for any } \alpha \neq 0,$$
which leads to our main result by providing the desired vanishing in the whole parameter range. One can  understand heuristically why by considering the contribution of the set where $|x_j-x_k| \lesssim \sqrt{\eps}$ and $|y_\ell|\lesssim \sqrt{\eps}$ for $\ell = 1 \ldots N$ to the right-hand side of~\eqref{eq:Hardy}. One has $|\bx_j - \bx_k| ^2 \lesssim \eps$ on this set. If the limiting 1D function does not vanish for $x_j\sim x_k$, then $|\Psi|^2 \propto U_\eps ^2 \propto \eps^{-N/2}.$ The total contribution (volume times typical value of the integrand) would thus be of order $\eps^{-1/2}$, much larger than the expected energies, of order unity after removal of the contribution of $N e_\eps$ as in~\eqref{eq:ener}. More precisely this leads us to (see Lemma~4.8 in~\cite{RouYan-23b})

\begin{lemma}
    Let 
\begin{multline}\label{psi1d}
    \psi (x_1,\dots,x_N) :=  \int_{\R^N} \d y_1\cdots \d y_N \\
      \left|\phi (\mathbf{x}_1,\dots, \mathbf{x}_N) \right| U_1^2 (y_1,\dots, y_N) \1_{|y_1|\leq 1/4} \ldots \1_{|y_N|\leq 1/4}  .
\end{multline}
and $K$ be a bounded open subset in $\mathbb{R}^{N-1}$. For a constant $\gamma \ge 0$, we define 
    \begin{equation*}\label{defKgamma}
        K_{\gamma} := \left\{ (x_2,x_3,\dots,x_N) \in K : | x_2-x_j| > \sqrt{\gamma}/2, \forall j \ge 3\right\}.    
    \end{equation*}
    Then when $\gamma\ge \varepsilon$, we have
    \begin{equation*}
     \int_{K_{\gamma}} \left| \psi (x_2,x_2,x_3,\dots,x_N) \right|^2 \d x_2 \cdots \d x_N 
    \le C_K \varepsilon^{\frac{s}{2}}\label{newEstm}
\end{equation*}
for some constants $C_K>0$  and $s \in (0,1)$ both independent of $\varepsilon$.
\end{lemma}

The above will, after passing to the limit, guarantee that $\psi_0$ vanishes when two arguments come close, the others being at least at a distance $\sqrt{\gamma}/2$. One can finally pass to the limit $\gamma \to 0$ to deduce that $\psi_0 (x_j=x_k) \equiv 0$ for any $k\neq j$. Hence indeed the energy bounds force the limiting 1D function to vanish upon particle encounters, as desired. 

%

\section{Conclusions}

We have studied 2D anyons of statistics parameter $\alpha$ in the magnetic gauge picture, i.e. seeing them as bosons in a varying magnetic field proportional to matter density. In a cold atoms context this corresponds to proposals made e.g. in~\cite{ValWesOhb-20}. We imposed a dimensional reduction by ways of a strongly anisotropic trap, as in recent cold-atoms experiments probing density-dependent gauge theories~\cite{ClaEtalChi-18,FroEtalTar-22,YaoZhaChi-21}. 


In the 1D limit we found that a suitable choice of gauge removes long-range magnetic interactions. Their only remnant is a hard-core condition upon particle encounters, leading to the Girardeau-Tonks model of 1D bosons for any $\alpha \neq 0.$ Non-trivial dependence on $\alpha$ might survive at sub-leading order, in which case it could be determined by perturbation theory around Girardeau's solution of the impenetrable 1D Bose gas.

\bigskip

\noindent\textbf{Acknowledgments.} 
Funding from the European Research Council (ERC) under the European Union's Horizon 2020 Research and Innovation Programme (Grant agreement CORFRONMAT No 758620) is gratefully acknowledged. We had interesting discussions with Douglas Lundholm, Michele Correggi, Per Moosavi and Nikolaj Thomas Zinner.

\end{document}